OpenStudy (anonymous):

Write the equation of a line with the slope, 3/2 which passes through the point (2,1) write the answer in slope-intercep form

OpenStudy (anonymous):

(y - 1) = (3/2)*(x - 2)

OpenStudy (anonymous):

First put in point slope formula which is y-(y)=slope(x-(x)) This is just simply stating the you minus y from the y coordinate= slope(x- x coordinate) Once you simplify the equation you will get in in slope-intercept form.

OpenStudy (anonymous):

So how do I write it in my answer? I tried the first answer and it's saying it's wrong....

OpenStudy (anonymous):

y-1 = (3/2)*(x - 2) y = (3/2)*x -3 + 1 y = (3/2)*x - 2

OpenStudy (anonymous):

Got it! Thank you! :)